A company is projected to generate free cash flows of $643 million per year for the next 3 years (FCFF1, FCFF2 and FCFF3). Thereafter, the cash flows are expected to grow at a 2.7% rate in perpetuity. The company's cost of capital is 8.1%. The company owes $179 million to lenders and has $59 million in cash. If it has 189 million shares outstanding, what is your estimate for its stock price? Round to one decimal place.
Answer #1

FCF1 = $643
FCF2 = $643
FCF3 = $643
FCF4 = $643 * (1 + 2.7%) = $660.361

Terminal value = FCF4 / (Ke - g) = $660.361 / (0.081 - 0.027) = $12,228.90741

Current total value of firm = ($643 / 1.081) + ($643 / 1.081^2) + ($643 / 1.081^3) + (12,228.91 / 1.081^3)
= $11,334.8729

Equity value = Current firm value + cash - debt
= $11,334.87 + $59 - $179
= $11,214.8729

Share price = Equity value / Number of shares
= $11,214.8729 / 189
= $59.3

Share price = $59.3